OpenStudy (anonymous):

Hello, What is the general solution of the following differential equation? (1 + (x^3/y)sin^2x)dx + (1/y)(x + (1/cos^2(2y))dy = 0; I used the integrating factor, but I'm stuck at the following step: u(x) [-x^3sin^2x + 1] = u'(x) [x + (1/cos^2(2y)] Thanks in advance!

OpenStudy (anonymous):

The original eq is: \[(1 + (x ^{3}/y)\sin ^{2} x)dx + (1/y)(x + (1/\cos ^{2}(2y))dy = 0\] I'm stuck at this step: \[u(x) [-xsin ^{2} x + 1] = u'(x) [x + (1/\cos ^{2}(2y)]\]